1. In class, we derived the Central Difference formula for f(x0) as f(x) = f(xo + h) – f(20 - h) +002). f(o)= 2h A similar

Five-Point Midpoint Formula • f(xo) = [f(xo – 2h) – 8f (xo – h) +8f (xo +h) - f (xo + 2h)] + with f(5) (5), (4.6) where liesPlease answer question #1 and write legibly - I have also equation 4.6 from the textbook. Thanks
Answer #1

about x = No is , We have to prove : = f(x) = [ 62-21) – 8f66-n) +8f69+h) - f(x + 1)] + 5* 4°() The taylor series formula fo* (20) * () * () * ( ب ) = (24+ بها یا د ) (5) - (تام ها (%0) * ( - ) = Ak) (4) () + ( : مواد (1) و (2) و (954 م2) - (+) 8 (1B - [fla)t on files to the ) + 8? 4 (2) + e pa lan) 1 sp Am] = f() [2-848-1]+ 1(») [-2h +8h +8 17:36] • 8h +8 ha = (x) *0:: 184 (2) - Apna jºlf.) - +7 -n) - 8 f(x-1) + 8 +(4+h) - (+dh) 4x30 where į is between 2-2h and to t2h . We have neglected( 2) A[(x - A)-3 (a + + f(a+b) - spon) + s (5) 以)